What is the probability that Mike spends at most 240 min on online gaming if his mean time spent on it is 270 min with a standard deviation time of 15 min?
1
Expert's answer
2020-10-27T18:42:27-0400

Let's assume that time spent is normally distributed.

"P(x<240)=P(z<\frac{240-270}{15})"

"=P(z<-2)=0.0228"